Accents Package version 1.0a

Javier Bezos-López^{∗}

August 3, 1997

1 Description

This package reimplements the accent commands \hat, \check, \tilde, etc., to allow multiple

accents with right placement. Also adds the accent \ring (x̊). So you can write ˆ ˆ ĥ with $\hat{\hat{\hat h}}$.^{1} The code is brief but slow, although in fast systems, like Intel 486, Motorola 68030 or RISC technology, is hardly felt. Also, a new command is defined: \accentset{acc-symbol}{nucl-symbol} which builds an

accent, as for instance \accentset{\star}{d} (^{⋆} d) or \accentset{\diamond}{h} (^{⋄}h). Accents

so defined behaves in quite similar way to real accents. Note that acc-symbol is always in \scriptscriptmode; hence, using \accentset in scripts won’t give the desired result. The command \underaccent{accent}{nucl-symbol} puts the accent under the symbol, but only “real” accents are allowed. For instance, \underaccent{\hat}{x} (x̂). \undertilde is the equivalent to \widetilde. Some points are to be noted:

• Slant of letters are automatically taken care of, even in \accentset.

• Placement of scripts after the symbol are right, as for instance in

     \hat{\hat{\hat f}}_1^1 ( ˆ ˆf̂_{1}^{1})

• The package does not provides alternative accents for fonts lacking them. If you want an accented \mathcal letter you had to write \hat{\hat{\mathcal{A}}} ( ˆÂ).

   • \mathbf{\hat{\hat h}} gets the bold accent; \hat{\hat{\mathbf{h}}} not.

• The reimplemetation of accents presumes standard encodings. (This point needs improvement.)

Two package options are provided:

single If you are interested only in \accentset. Accents are not reimplemented.

nonscript If you intend to use multiple accents in text and display modes only.

Macros are speeded up dramaticaly with both options; if your system is slow, I commend using the options in drafts and removing them for the final print.

∗E-mail: jbezos@mx3.redestb.es

1The algorithm of the American Mathematical Society, at least in the early versions, misplaces the accents over the h.

2 Implementation

For the moment the code is not described, but I give a few hints: • The algorithm is (almost) recursive.

• Definitions are global so that the values in the last loop will be taken.

• Accents are built over an empty box with the width of the nucleous. The nucleous is put after, in order attach the scripts.